Suppose that 387 people will travel on a shuttle with room for 420 people. Each compartment seats 14 people before the next compartment opens. How many people will ride in the last compartment that opened?
what I did was I tried 387/14 and got about 27 seats for the full 14 compartment seats and I think the answer is 9 but not to sure
: Suppose that 387 people will travel on a shuttle with room for 420 people. Each compartment seats 14 people before the next compartment opens. How many people will ride in the last compartment that opened?
what I did was I tried 387/14 and got about 27 seats for the full 14 compartment seats and I think the answer is 9 but not to sure

Answer by vleith(1174)

387/14 = 27.642
So you have 27 full compartments and one that has some people in it, but not yet full
27*14 = 378 people are contained in the 27 full compartments.
that leaves 387-378 = 9 in the one that is not yet full
Also (14* 0.642) = 9 - so that 'remainder' is 9 in both cases